OpenStudy (anonymous):

Solve the following system of equations by elimination: 6x + 3y= 15 2x - 3y = 9.

OpenStudy (anonymous):

Add the terms 6x + 3y= 15 2x - 3y = 9 6x+2x= 8x 3y + (-3y) =0 15 +9=24 therefore 8x +0 = 24 Solve for x then substitute back into the original equation to find y